OpenStudy (anonymous):

Please help me simplify a quadratic equation. (x^2 + 2x +1) + (y^2 + 4y +4) = 25

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you seem to have two variables there, but each term in parentheses is a perfect square

OpenStudy (misty1212):

you can rewrite it as \[(x+1)^2+(y+2)^2=25\] so it is the equation for a circle

OpenStudy (anonymous):

Yes, that is what we were supposed to do. It is to find the midpoint of a circle along with the radius

OpenStudy (misty1212):

it is not a "quadratic equation" nor can you "simplify" it, but you can write it in the standard from for a circle

OpenStudy (anonymous):

Terribly sorry, my internet likes to go out on me

OpenStudy (anonymous):

How would I write it in standard form ?

OpenStudy (misty1212):

it is not the internet, is is OS today

OpenStudy (anonymous):

Oh okay so how would I write it in standard form like you said?

OpenStudy (misty1212):

i did already

OpenStudy (anonymous):

Oh okay, I see

OpenStudy (misty1212):

\[x^2+2x+1=(x+1)^2\] and \[y^2+4y+4=(y+2)^2\]

OpenStudy (misty1212):

you get \[(x+1)^2+(y+2)^2=25\] and from that you can read off the center and the radius

OpenStudy (misty1212):

you got that or no?

OpenStudy (anonymous):

I used an online calculator to try this and ended up with "The center is (-1,-2) and the radius is 5." Would this be correct? Or would my points have to be positive

OpenStudy (anonymous):

And yes, I understand now

OpenStudy (misty1212):

that is correct form is \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

Okay, Thank you for all of your help :)